OpenStudy (anonymous):

quadratic formula to solve x2 + 8x = –2

OpenStudy (anonymous):

X^2+8X+2=0 X=-8+ ROOT(56) ------------ 2 OR X=-8 - ROOT(56) ------------ 2

OpenStudy (ash2326):

ax^2+bx+c=0 \[ x=(-b+\sqrt(b^2-4ac))/2a\] \[ x=(-b-\sqrt(b^2-4ac))/2a\]

OpenStudy (ash2326):

here a=1 b=8 c=2 b^2-4ac=56 so \[ x=(-8+\sqrt(56))/2\] or \[ x=(-4+\sqrt(14))\] and \[ x=(-4-\sqrt(14))\]

OpenStudy (anonymous):

Thnk you<3